three times the complement of an angle less the suppliment of an angle is 10 decrees. find the suppliment
Answer by stanbon(75887)

three times the complement of an angle less the suppliment of an angle is 10 decrees. find the supplement
-------
angle: x
complement: 90-x
supplement: 180-x
-------------------------
Equation:
3(90-x)-(180-x) = 10
270-3x-180+x = 10
-2x +90 = 10
-2x = -80
x = 40
------
supplement: 180-x = 180-40 = 140 degrees
=============================================
Cheers,
Stan H.
==================